HUBBLE SPACE TELESCOPE DAILY REPORT #2619 PERIOD COVERED: 0000Z (UTC) 05/08/00 - 0000Z (UTC) 05/09/00 Daily Status Report as of 130/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD/MA1 8040 (Dynamics and Energetics of Stellar Flares) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to perform Mode 1.3 echelle spectra will be used to explore the dynamics and energetics of flares on M dwarfs. The observations completed with no reported problems. 1.2 Completed Two Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working. The proposal completed with no reported problems. 1.3 Completed Three Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.4 Completed STIS/CCD/MA1 8241 (A Snapshot Survey of Interstellar Absorption Lines) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make high resolution Echelle spectrographic observations of the standard ultraviolet star HD122879 as part of a snapshot survey of interstellar absorption lines. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.6 Completed Three Sets of WF/PC-2 8346 (Cosmological Parameters from Type Ia Supernovae at High Redshift) The WF/PC-2 was used to directly measure the cosmological parameters, Omega, Lambda, and thus the curvature, Omega_k, using Type Ia supernovae {SNe Ia} as calibrated standard candles. Observations of SNe Ia spanning a large redshift range are key to decoupling Omega and Lambda. The observations completed nominally. 1.7 Completed Four Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.8 Completed WF/PC-2 8118 (Snapshot Survey of Galactic Globular Clusters0 The WF/PC-2 was used to make observations of crowded centers of globular clusters {GCs} which have produced intriguing discoveries of new phenomena, most of which are not understood theoretically. This is a continuation of a Cycle 7 snapshot proposal, to extend such work to a much larger portion of the Galactic GC system. The observations completed nominally, with no reported problems. 1.9 Completed Four Sets of WF/PC-2 8442 (Cycle 8 Standard Darks) The WF/PC-2 was used to obtain dark calibration dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported anomalies. 1.10 Completed Two Sets of STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.11 Completed STIS/CCD 8128 (Stalking FLIERs in Planetary Nebualae: a STIS Study) The Space Telescope Imaging Spectrograph (CCD) was used to observe FLIERs, which are enigmatic pairs of small, low-ionization knots located on opposite sides of the central star of many PNe. Their symmetric locations, equal, opposite and supersonic Doppler shifts {Mach 3-5}, and highly enriched N/O ratios suggest that they are peculiar spitballs ejected recently from the stellar interior. The observations completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 7 Scheduled Re-acquisitions: 5 Successful: 5 2.2 FHST Updates: Scheduled: 15 Successful: 15 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared once. A TTR was written for a 23-minute period of non-recoverable real-time data loss that began at 129/2130Z due to a WSGT upgrade in progress. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW